what is the answer to this problem?
5u+6-3(4u-1)=3(u-1)

Answer by lenny460(764)

5u + 6 - 3(4u - 1) = 3(u - 1)
-3 * 4u = -12u
-3 * -1 = 3
3 * u = 3u
3 * -1 = -3

5u + 6 - 12u + 3 = 3u - 3
Combine like terms:
5u - 12u = -7u
6 + 3 = 9

-7u + 9 = 3u - 3
Add 3 to each side of the equation
-7u + 9 + 3 = 3u - 3 + 3
Combine like terms:
9 + 3 = 12
-3 + 3 = 0
-7u + 12 = 3u
Add 7u to each side of the equation
-7u + 12 + 7u = 3u + 7u
Combine like terms:
-7u + 7u = 0
3u + 7u = 10u
12 = 10u or
10u = 12
Divide each side by 10
10u/10 = 12/10
u = 12/10 = 6/5 = 1 1/5
The Answer:
u = 6/5 = 1 1/5


obuong3@aol.com